OpenStudy (anonymous):

Eliminate the parameter. x = t - 3, y = t2 + 5

OpenStudy (helder_edwin):

in the first equation solve for t, then replace it into the second equation.

OpenStudy (anonymous):

y = x2 + 14?

OpenStudy (helder_edwin):

NO. from the first equation \[\large t=x+3 \] so \[\large y=t^2+5=(x+3)^2+5 \]

OpenStudy (anonymous):

This has to be it: y = x2 + 6x + 14

OpenStudy (helder_edwin):

yes. but there's (usually) no need to expand the square: u lose information.

OpenStudy (helder_edwin):

(i mean u hide it, u don't lose it, it is still there)